OpenStudy (anonymous):

What transformations change the graph of (f)x to the graph of g(x)? f(x) = x² ; g(x) = (x + 7)² + 9 The graph of g(x) is the graph of f(x) translated to the down 7 units and right 9 units. The graph of g(x) is the graph of f(x) translated to the up 7 units and left 9 units. The graph of g(x) is the graph of f(x) translated to the right 7 units and down 9 units. The graph of g(x) is the graph of f(x) translated to the left 7 units and up 9 units.

OpenStudy (anonymous):

the parent function is x^2, it seems that x^2 is becoming (x+7)^2, this transformation is like F(x)=x^2 => F(X+7)=(x+7)^2 the result of this is shifting the graphs origin 7 units left The addition of 9 is an addition outside the parent function which shifts it up or down. therefore g(x) is a linear transform of 7 units left and 9 units up